having a problem with teh following question? A shirt is marked 20% off and the sale price is $32.00. What was the original price?

Answer by filboi(35)

You may use two solutions.
Solution 1:
Original Price - 20% of Original Price = Discounted Price
x - 0.20x = 32
0.80x = 32
x = 40

Solution 2:
You know the 32 is only 80% of the original price. So use, ratio & proportion.
32/80% = x/100%
x = 40
So the answer is $40.